pter 5 Parrino, Fundamentals of Corporate Finance, 4e Grace Period: 9 days left l Register PRINTER VERSION 4 BACK NEXT MENT CES Problem 5.26 You have just inherited $550,000. You plan to save this money and continue to live off the money that you are earning in your current job. If you can invest the money in a bond that pays 4,18 percent interest annually, how long will it be before your inheritance is worth $1 intermediate calculations e.g. 8.72.) million? (I you solve this problem with aigebra round to 5 decimal places, in all cases round your final answer to 2 decimal places 5.03 Inheritance will be worth $1 million in years 5.16 5.19 (Excel 5.22 Question Attempts: O of 2 used SAVE FOR LATER SUBMIT ANSWER Videol e靡 ENG 1046 PM 1/17/2019
Answer #1
Amount of investment            550,000
Interest 4.18%
Amount required        1,000,000
Time N
Amount= Investment * (1+ Interest)^Time
1000000= 550000*(1+4.18%)^N
=1000000/550000 = 1.0418^N
1.818181818 = 1.0418^N
Log 1.8181= N * Log 1.0418
0.25963731= N*0.01778435
N= 0.25963731/0.01778435
N=                14.60
So in 14.60 years the amount will become 1 million